### v2

# Hamiltonian stochastic Lie systems and applications

Abstract
This paper provides a practical approach to stochastic Lie systems, i.e. stochastic differential equations whose general solutions can be written as a function depending only on a generic family of particular solutions and some constants related to initial conditions. We correct the stochastic Lie theorem characterising stochastic Lie systems proving that, contrary to previous claims, it retains its classical form in the Stratonovich approach. Meanwhile, we show that the form of stochastic Lie systems may significantly differ from the classical one in the Itô formalism. New generalisations of stochastic Lie systems, like the so-called stochastic foliated Lie systems, are devised. Subsequently, we focus on stochastic Lie systems that are Hamiltonian systems relative to different geometric structures. Special attention is paid to the symplectic case. We study their stability properties and devise the basics of a stochastic energy-momentum method. A stochastic Poisson coalgebra method is developed to derive superposition rules for Hamiltonian stochastic Lie systems. Potential applications of our results are shown for biological stochastic models, stochastic oscillators, stochastic Lotka–Volterra systems, Palomba–Goodwin models, among others. Our findings complement previous approaches by using stochastic differential equations instead of deterministic equations designed to grasp some of the features of models of stochastic nature.
Keywords: energy-momentum method, Hamiltonian stochastic Lie system, Poisson coalgebra, stability, stochastic Lie system, superposition rule, Stratonovich formalism.
MSC 2020: 60H10, 34A26 (Primary), 37N25, 53Z05 (Secondary).
1 Introduction
In its most classical definition, a superposition rule is a $t$-independent function that describes the general solution of a $t$-dependent system of first-order ordinary differential equations (ODEs) in normal form, a so-called Lie system, via a generic family of its particular solutions and a set of constants related to initial conditions [9, 17, 18, 56]. Superposition rules are used, for instance, in approximate and numerical methods, as they are available for Lie systems whose exact general solutions in explicit form are unknown [20, 51, 56]. In particular, the knowledge of a particular finite set of exact and/or approximate solutions of a Lie system permits one to study its general solution via superposition rules [56].
The superposition rule concept traces its origins back to Sophus Lie’s pioneering and celebrated book [44], redacted and edited by Georg Scheffers. In that work, Lie stated the nowadays called Lie theorem characterising Lie systems. Previous to that, Lie briefly described his Lie theorem in [43] without a proof as a criticism to some previous works on superposition rules by Vessiot, Guldberg, Köningsberger, and other researchers (see [23, 34, 54] and references therein). Lie stated that the results on superposition rules for differential equations on $\mathbb{R}$ devised by previous authors were a simple application of his theory on infinitesimal transformation groups. Moreover, [44] laid down the foundations for the theory of Lie systems. The Lie theorem is also called Lie–Scheffers theorem, as Scheffers took part in writing Lie’s work [44], or Lie superposition theorem [10]. Previous remarks suggest one to say ‘Lie theorem’ rather than ‘Lie–Scheffers theorem’ and to use the denomination ‘Lie system’ instead of ‘Lie–Scheffers system’. Since Vessiot devised important contributions to the theory of Lie systems [20, 23], Lie–Vessiot system is also a good denomination for a Lie system. On the contrary, Scheffers never worked on Lie systems by himself [23].
The Lie theorem states that a $t$-dependent system of ODEs in normal form admits a superposition rule if and only if it describes the integral curves of a $t$-dependent family of vector fields that can be considered as a curve in a finite-dimensional Lie algebra of vector fields, a Vessiot–Guldberg Lie algebra (see [10, 17, 18] for modern approaches and further details).
Lie systems have been thoroughly studied due to their widespread occurrence in physics and mathematics (see [20, 23] counting more than 300 references on Lie systems and related topics). Towards the 1980s, Winternitz and his colleagues at the Centre de Recherches Mathématiques of the University of Montreal conducted an extensive study on Lie systems. Subsequently, many other authors followed their work [23]. There has been a bast effort to generalise Lie systems to more general situations: $t$-dependent Schrödinger equations [19, 20], quasi–Lie systems [16] and quasi–Lie families [14], foliated Lie systems [17, 21], discrete differential equations [50], stochastic differential equations [39], superdifferential equations [8], et cetera [13]. In this work, we are mainly concerned with the extension of superposition rules and Lie systems to the realm of stochastic differential equations [39].
Stochastic differential equations may describe phenomena that deterministic differential equations can not [2]. For example, the probability of disease extinction or outbreak, the quasi-stationary probability distribution, the final size distribution, and the expected duration of an epidemic are features that cannot effectively be modelled by deterministic methods [2]. These and other reasons motivate the huge interest in studying stochastic differential equations. Nevertheless, some deterministic differential equations can grasp certain interesting characteristics of models even easier than stochastic differential equations [12] and stochastic differential equations offer an alternative view that can be complementary [33].
It is interesting to extend superposition rules to stochastic differential equations. The work [39] extends the superposition rule notion to a class of systems of stochastic first-order ordinary differential equations, called stochastic Lie systems. The authors use a Stratonovich approach, as this makes the approach similar to the deterministic theory of Lie systems. Moreover, [39] presents an interesting and very precise presentation of certain local results about stochastic Lie systems and, as a byproduct, it also explains many technical results on standard Lie systems, which use to be absent in the literature [17, 18, 23]. Despite its mathematical interest, many of the technical details given in [39] are frequently omitted as they have, in general, not many practical applications. As noted in [10, 23], standard works on Lie systems, even theoretically ones, are essentially interested in local aspects and generic points, which makes them skip many technical details analysed in [39]. Notwithstanding, there is mathematical interest in global and local technical aspects, as illustrated by the work [10] for global superposition rules and [39]. It is worth stressing that stochastic Lie systems were found to have applications in the description of Brownian motions, economical models like the Black–Scholes theory of derivatives prices, and so on [39]. The potential interest of stochastic Lie systems in epidemic models was very briefly mentioned in [27], without, as far as we know, any further development.
In [39], a stochastic Lie theorem characterising Stratonovich stochastic first-order ordinary differential equations admitting a superposition rule was devised, but it contains a mistake that changes its meaning and potential applications. More exactly, the direct part of the stochastic Lie theorem in [39] states that a stochastic Lie system may admit, locally, a superposition rule if its associated Stratonovich operator is related to a family of vector fields spanning an involutive distribution. Our present work proves that the vector fields must additionally span a finite-dimensional real Lie algebra of vector fields, which is a much stronger condition appearing already in the classical Lie–Scheffers theorem [17, 18, 44]. We also determine the precise point of the mistake in [39], we correct the statement of the stochastic Lie theorem, and we provide a counterexample, based on SIS epidemic models, to illustrate when the stochastic Lie theorem in [39] fails. It is worth stressing that the differences between the wrong and the correct versions of the stochastic Lie theorem have no impact in the applications carefully studied in [39].
Moreover, our work presents a concise introduction to stochastic Lie systems, aiming at providing a practical approach while avoiding technical details that are not necessary for general purposes. In this sense, it simplifies the elegant and mostly rigorous mathematical treatment in [39] by using standard assumptions in mathematical constructions. For instance, we focus on local results at generic points, which significantly simplifies previous needed techniques.
Our correction of the stochastic Lie theorem shows that the stochastic Lie theorem has no exclusive stochastic features in the Stratonovich approach: it retains the conditions of the classical Lie theorem. Meanwhile, it is stressed that the conditions for a stochastic differential equation in the Itô form [53] to become a stochastic Lie system do not follow the standard form expected from the deterministic Lie theorem [17, 18, 20]. This is very important practically, as many relevant stochastic differential equations are given in the Itô form and they have to be translated into the Stratonovich approach [52] to apply the methods of our work and [39]. In this respect, the relation between the Itô and the Stratonovich approaches is reviewed and some examples with potential applications are provided in this work. It is worth noting that stochastic differential equations in the Itô form may seem to be stochastic Lie systems, but they are not. This occurs in certain SIS epidemiological models [32].
Apart from introducing stochastic Lie systems, this paper suggests the applicability of various generalisations of Lie systems [20] to the stochastic domain. This unveils a new vast realm of potential applications in physics, mathematical, and biological models, offering promising avenues for further exploration. One example is the extension of the theory of foliated Lie systems [21] to the realm of stochastic differential equations. In fact, we suggest that, in analogy with the deterministic case, this can be the case when studying certain problems of our here devised stochastic energy-momentum method [25].
By the stochastic Lie theorem, every stochastic Lie theorem admitting $\ell$ independent random variables is related to a family of $\ell$ vector fields (if the time is considered as a special ‘deterministic’ random variable) parametrised by random variables. All of them can be understood as linear combinations of elements of a finite-dimensional Lie algebra of vector fields (a so-called Vessiot–Guldberg Lie algebra) with coefficients depending on the random variables. More particularly, we study stochastic Lie systems admitting a Vessiot–Guldberg Lie algebra of Hamiltonian vector fields relative to some compatible differential geometric structure: the Hamiltonian stochastic Lie systems. In particular, we mainly focus on Hamiltonian stochastic Lie systems relative to symplectic forms, although our theory is easy generalisable to other geometric structures and stochastic Lie systems. In this context, the coalgebra method is extended to Hamiltonian stochastic Lie systems to derive superposition rules. This provides an extension to the stochastic realm of the theory of Hamiltonian Lie systems and their generalisations [23].
Our results are illustrated with many new examples of stochastic Lie systems. In particular, we study SIS models [32]. SIS models are epidemiological models that assume that individuals do not acquire immunity after infection. They concern two variables/compartments: $S$ representing susceptible individuals and $I$ representing infected individuals in a large population of size $N$ where a single disease is spreading. SIS systems are usually employed in the literature in a deterministic manner [27]. This can be used to describe some of their features, but not all, as some are of pure stochastic nature. Our models can also be used to study stochastic models appearing in Lotka–Volterra models, Palomba–Goodwin models [49], stochastic oscillators [35, 36], etc. In fact, stochastic Lie systems in general and Hamiltonian stochastic Lie systems, in particular, seem to have a quite large field of potential applications.
We are also concerned with a theory of stability for Hamiltonian stochastic systems and, in particular, Hamiltonian stochastic Lie systems. Our study is specially concerned with linear ones, which appear as an approximation of nonlinear ones and they share some stability properties with them [4]. Moreover, the basis for a energy-momentum method [46] for Hamiltonian stochastic differential equations is laid down by using some results in [40]. In particular, this allows one to study the relative equilibrium points of Hamiltonian systems, i.e. points where the dynamics is generated by Hamiltonian symmetries of the system under study. A characterisation (see Theorem 6.5) of the relative equilibrium points for stochastic Hamiltonian systems in terms of critical points of its Hamiltonian functions is developed. As a byproduct, strong constants of motion [40] for Hamiltonian stochastic Lie systems are briefly studied and illustrated with examples.
The structure of the paper goes as follows. Section 2 is a brief introduction to stochastic differential equations, stochastic Lie systems, and many other notions to be used in this paper. It also shows the difference of the form of stochastic Lie systems in the Stratonovich and the Itô approaches. Section 3 is concerned with superposition rules for stochastic Lie systems and it reviews and corrects the previous version of the stochastic Lie system. Section 4 deals with Hamiltonian stochastic Lie systems. In particular, we define the newly proposed stochastic foliated Lie systems, and the Hamiltonian counterparts of stochastic Lie systems. A theory of stability for stochastic Lie systems is given in Section 5. A relative equilibrium notion is presented and studied in Section 6 and a stochastic version of a classical result characterising relative equilibrium points is presented. Meanwhile, Section 7 develops the Poisson coalgebra method for Hamiltonian stochastic Lie systems. Finally, Section 8 presents our conclusions and future work.
2 Stochastic differential equations and stochastic Lie systems
This section provides a concise introduction to stochastic differential equations and stochastic Lie systems. A detailed survey on the theory of stochastic differential equations can be found in [4, 26], while the theory of the stochastic Lie systems was elaborated for the first time in [39], which offers a precise formulation of the theory. To avoid technical details, we will assume objects to be smooth, locally defined, and problems at generic points satisfying very mild conditions. Following the classical Lie systems theory [17, 20, 44], we here provide a stochastic Lie system definition not based on the superposition rule notion.
In a nutshell, a $t$-dependent system of first-order ordinary differential equations on an $n$-dimensional manifold $M$ of the form
$\frac{\mathrm{d}\Gamma^{i}}{\mathrm{d}t}=X^{i}(t,\Gamma)\,,\qquad i=1,\ldots,n\,,\qquad\forall t\in\mathbb{R}\,,\qquad\forall\Gamma\in M\,,$ (2.1)
for certain functions $X^{1},\dotsc,X^{n}\in\mathscr{C}^{\infty}(\mathbb{R}\times M)$, is deterministic in the sense that an initial condition in $M$ at a time $t_{0}$ establishes, under mild conditions on the functions $X^{1},\ldots,X^{n}$, a unique solution giving a position in $M$ for every $t\in\mathbb{R}$. Then, (2.1) may be modified by considering that its form can also depend on certain ‘stochastic processes’ $B^{1},\ldots,B^{r}$, namely a series of $t$-dependent random variables satisfying certain appropriate conditions to be set hereafter in detail. This fact is shown by the expression
$\delta\Gamma^{i}=X^{i}_{1}(B,\Gamma)\delta t+\sum_{\alpha=2}^{\ell}X^{i}_{\alpha}(B,\Gamma)\circ\delta B^{\alpha}\,,\qquad i=1,\ldots,n\,,$ (2.2)
for functions $X^{i}_{1},\ldots,X_{\ell}^{i}\in\mathscr{C}^{\infty}(\mathbb{R}^{\ell}\times M)$, with $B=(t,B^{2},\ldots,B^{\ell})$, $\Gamma=(\Gamma^{1},\ldots,\Gamma^{n})$, and $i=1,\ldots,n$. The symbol $\circ$ has been used to indicate that (2.2) is understood in the so-called Stratonovich interpretation, to be briefly explained afterwords. More precisely, $(\Omega,\mathcal{F},P)$ is a probability space, where $\Omega$ is a manifold, $\mathcal{F}$ is a $\sigma$-algebra of subsets of $\Omega$, and $P:\mathcal{F}\rightarrow[0,1]$ is a probability function on $\mathcal{F}$. Each $B_{\alpha}:\mathbb{R}_{+}\times\Omega\rightarrow\mathbb{R}$ is a semi-martingale for $\alpha=1,\ldots,r$. Semi-martingales are good integrators relative to the Itô or the Stratonovich integrals due to its properties. In short, a martingale is a sequence of stochastic processes such that, at a particular time, the conditional expectation of the next value is equal to the present value, independently of all previous values.
A stochastic differential equation is then an expression on a manifold $M$ of the form [^1] Since $\mathfrak{S}$ stands for a Stratonovich operator, we will sometimes write $\mathfrak{S}(X,\Gamma)\delta X$ instead of $\mathfrak{S}(X,\Gamma)\circ\delta X$ to simplify the notation when it is clear what we mean from context.
$\delta\Gamma=\mathfrak{S}(B,\Gamma)\circ\delta B\,,$ (2.3)
where $B:\mathbb{R}_{+}\times\Omega\rightarrow\mathbb{R}^{\ell}$ is an $\mathbb{R}^{\ell}$-valued semi-martingale and $\mathfrak{S}(B,\Gamma):\mathrm{T}_{B}\mathbb{R}^{\ell}\to\mathrm{T}_{\Gamma}M$, with $(B,\Gamma)\in\mathbb{R}^{\ell}\times M$, describes a Stratonovich operator. Every basis in $\mathrm{T}^{\ast}\mathbb{R}^{\ell}$ allows one to decompose $\mathfrak{S}(B,\Gamma)$ into $\ell$-components $(\mathfrak{S}_{1}(B,\Gamma),\ldots,\mathfrak{S}_{\ell}(B,\Gamma))$ in the chosen basis, which will be frequently employed hereafter. Every particular solution to (2.3) is also a semi-martingale $\Gamma:\mathbb{R}_{+}\times\Omega\rightarrow M$. Moreover, we say that a particular solution has initial condition $\Gamma_{0}\in M$ when $\Gamma(0,\omega_{0})=\Gamma_{0}$ for every $\omega_{0}\in\Omega$ with probability one. Note that the standard time can be considered as a random variable $t:(t,\omega_{0})\in\mathbb{R}_{+}\times\Omega\mapsto t\in\mathbb{R}$, whose value is independent of $\Omega$ and to be included as the first component of $B$. In practice, $B$ is related to Brownian motions, also called Wiener processes, càdlág martingales, Itô processes of the form $\delta X=\sigma\delta W+\mu\delta t$ for a Brownian motion $W$ and adapted processes $\sigma,\mu$, Lévy processes, etc. It is worth noting that semimartingales form the largest class of processes for which the Itô integral can be defined. On the other hand, white noises are not semi-martingales.
The solution to (2.3) is of the form
$\Gamma(t)-\Gamma(0)=\int_{0}^{t}\mathfrak{S}_{1}(B,\Gamma)\delta t+\sum_{\beta=2}^{\ell}\int_{0}^{t}\mathfrak{S}_{\beta}(B,\Gamma)\circ\delta B^{\beta}_{t},$ (2.4)
where the integrals appearing above are Stratonovich integrals. It is relevant to understand that stochastic differential equations can still be understood in the so-called Itô way, in which the general solution is also of the form (2.4), but the integrals are assumed to be Itô integrals, which is different. If one approach or the other is used, depends on the applications to be developed.
Unless otherwise stated, every stochastic differential equation (2.3) is assumed to be in the Stratonovich sense. This is motivated by the Malliavin transfer principle [42, 45], which suggests that the obtained theory will retain the standard differential theory, although it may not be always this way. In other words, Malliavin transfer principle states that “A formula which is true in the deterministic context and which has a meaning via Stratonovitch stochastic calculus, is still valid, but only almost surely” [42].
Despite this general idea, stochastic differential equations appear in the Itô sense in many applications [28, 48, 55],. Hence, a manner to deal with such stochastic differential equations will be studied in this work. Moreover, the analysis of such stochastic differential equations show some differences appearing in the stochastic formalism with respect to the deterministic theory of Lie systems, which enriches the theory.
Although every Stratonovich stochastic differential equation is equivalent to another Itô stochastic differential equation, the form of both is different. Hence, one has to take care of the method employed to study a stochastic differential equation [29]. More exactly, assume that (2.2) has coefficients that do not depend on ${B}^{2},\ldots,{B}^{\ell}$, then the Itô differential equation admitting the same solutions reads (see [11, p. 137] for details)
$\delta\Gamma^{i}=\left(\mathfrak{S}^{i}_{1}(t,\Gamma)+\frac{1}{2}\sum_{\beta=2}^{\ell}\sum_{j=1}^{n}\frac{\partial\mathfrak{S}_{\beta}^{i}}{\partial\Gamma^{j}}(t,\Gamma)\mathfrak{S}_{\beta}^{j}(t,\Gamma)\right)\delta t+\sum_{\beta=2}^{\ell}\mathfrak{S}^{i}_{\beta}(t,\Gamma)\delta{B}^{\beta}_{t},\qquad i=1,\ldots,n.$ (2.5)
Note that we have dropped the $\circ$ sign in the previous expression. It is worth stressing that the definition of a stochastic Lie system to be given soon relies on the new term appearing multiplying $\delta t$ in (2.5), which is called the drift term [11]. Moreover, the transformation from the Stratonovich to the Itô form does not change the coefficients with the $\delta{B}^{\beta}_{t}$ for $\beta=2,\dotsc,\ell$.
Example 2.1.
Let us consider a stochastic differential equation induced by a semi-martingale $B:(t,z)\in\mathbb{R}_{+}\times\Omega\mapsto(t,\mathcal{B})\in\mathbb{R}^{2}$ consisting of two variables (a deterministic one, $t$, which can be understood as a particular type of stochastic variable, and a purely stochastic one, $\mathcal{B}$, describing a Brownian motion) of the Itô form
$\delta I=I(\beta N-\mu-\gamma-\beta I)\delta t+I\sigma(N-I)\delta\mathcal{B}\,,$ (2.6)
where $N$ is a constant describing the total population of a SIS epidemiological system [32]. Let us assume that $\sigma=\sigma(t)$. This model describes the dissemination of a single communicable disease in a susceptible population of size $N$ (see [32]).
The Stratonovich stochastic differential equation related to (2.6) takes the form
$\delta I=\left(-\sigma^{2}(t)I^{3}+\left(-\beta+\frac{3N\sigma^{2}(t)}{2}\right)I^{2}+I\left(N\beta-\gamma-\mu-\frac{\sigma^{2}(t)N^{2}}{2}\right)\right)\delta t+\sigma(t)I(N-I)\circ\delta\mathcal{B}\,.$ (2.7)
This induces a Stratonovich operator between $\mathbb{R}^{2}$ and $\mathbb{R}$ such that, for each $t,\mathcal{B},I$, one has an operator
$\mathfrak{S}(t,\mathcal{B},I):(\delta t,\delta\mathcal{B})\in\mathrm{T}_{(t,\mathcal{B})}\mathbb{R}^{2}\longmapsto\mathfrak{S}_{1}(t,I)\delta t+\mathfrak{S}_{2}(t,I)\delta\mathcal{B}\in\mathrm{T}_{I}\mathbb{R}\,$
for
$\begin{gathered}\mathfrak{S}_{1}(t,I)=-\sigma^{2}(t)I^{3}+\left(-\beta+\frac{3N\sigma^{2}(t)}{2}\right)I^{2}+I\left(N\beta-\gamma-\mu-\frac{\sigma^{2}(t)N^{2}}{2}\right),\\ \mathfrak{S}_{2}(t,I)=\sigma(t)I(N-I).\end{gathered}$
Solutions to (2.7) are given by semi-martingales of the form $\Gamma:(t,\omega_{0})\in\mathbb{R}_{+}\times\Omega\mapsto I(t,\omega_{0})\in\mathbb{R}$ determined by an initial condition described by a random variable $\Gamma_{0}:\Omega\rightarrow\mathbb{R}$ such that $\Gamma_{0}(\omega_{0})=1$ for every $\omega_{0}\in\Omega$. $\triangle$
Recall that the theoretical utilisation of Stratonovich stochastic equations is justified by Malliavin’s Transfer Principle [45], which states that the results from the theory of ordinary differential equations remain applicable, in an analogous way, to stochastic differential equations in Stratonovich form. This principle is just a general guideline without a proof, which implies that it must be used just as a general suggestion. Nevertheless, as shown in Example 2.1, the relation to the Itô approach has to be considered too for analysing applications.
Definition 2.2.
A stochastic Lie system is a stochastic differential equation on a manifold $M$ of the form $\delta\Gamma=\mathfrak{S}(B,\Gamma)\circ\delta B$ such that $B:\mathbb{R}_{+}\times\Omega\rightarrow\mathbb{R}^{\ell}$ is a semi-martingale and $\mathfrak{S}$ is a Stratonovich operator such that
$\mathfrak{S}(B,\Gamma)=\left(\sum_{\alpha=1}^{r}b_{1}^{\alpha}(B)X_{\alpha}(\Gamma),\ldots,\sum_{\alpha=1}^{r}b_{\ell}^{\alpha}(B)X_{\alpha}(\Gamma)\right),\qquad\forall\Gamma\in M,\quad\forall B\in\mathbb{R}^{\ell},$ (2.8)
for family of functions $b_{a}^{\alpha}:B\in\mathbb{R}^{\ell}\mapsto b_{a}^{\alpha}(B)\in\mathbb{R}$, with $\alpha=1,\ldots,r$ and $a=1,\ldots,\ell$, and a certain $r$-dimensional Lie algebra of vector fields on $M$ spanned by $X_{1},\ldots,X_{r}$. We call the Lie algebra $V=\langle X_{1},\ldots,X_{r}\rangle$ a Vessiot–Guldberg Lie algebra of the stochastic Lie system (2.8).
Note that the Stratonovich operator $\mathfrak{S}$ of a stochastic Lie system can be considered as an $\ell$-element family consisting of $\mathbb{R}^{\ell}$-vector fields on $M$. Each element of the family can be understood as a section of the trivial bundle ×$\mathbb{R}^{\ell}\times V\rightarrow\mathbb{R}^{\ell}$, and conversely.
Example 2.3.
Let us consider a damped harmonic oscillator on $\mathrm{T}\mathbb{R}$ with adapted coordinates $\{x,y\}$ and with semi-martingale $W_{1}$ in Itô form (cf. [35, 36, 47]) given by
$\delta\begin{pmatrix}x\\ y\end{pmatrix}=\begin{pmatrix}0&1\\ -\omega_{f}^{2}&-k\end{pmatrix}\begin{pmatrix}x\\ y\end{pmatrix}\delta t+\begin{pmatrix}0&0\\ 0&-\sigma\end{pmatrix}\begin{pmatrix}x\\ y\end{pmatrix}\delta W_{1}\,.$ (2.9)
More specifically, $W_{1}$ is a one-dimensional Wiener process, which is a type of semi-martingale, while $\sigma$ is a parameter, e.g. quantifying noise. The parameters $\omega_{f}$ and $k$ are the usual parameters relative to the standard deterministic model for a dissipative harmonic oscillator
$\ddot{x}+\omega_{f}^{2}x+k\dot{x}=0\,,$
where $\omega_{f}$ is the elastic constant of the oscillator and $k$ is the friction coefficient. Let us see how this model can be considered as a linear stochastic Lie system.
Consider the vector fields on $\mathrm{T}\mathbb{R}$ given by
$X_{11}=x\frac{\partial}{\partial x}\,,\qquad X_{12}=y\frac{\partial}{\partial x}\,,\qquad X_{21}=x\frac{\partial}{\partial y}\,,\qquad X_{22}=y\frac{\partial}{\partial y}\,.$ (2.10)
These vector fields span a Lie algebra of vector fields $V_{D}$ isomorphic to the general Lie algebra, $\mathfrak{gl}(2,\mathbb{R})$, of $2\times 2$ matrices with real coefficients. In fact, the commutation relations in the basis (2.10) read
$[X_{11},X_{12}]=-X_{12}\,,$ $[X_{11},X_{21}]=X_{21}\,,$ $[X_{11},X_{22}]=0\,,$
$[X_{12},X_{21}]=X_{22}-X_{11}\,,$ $[X_{12},X_{22}]=-X_{12}\,,$ $[X_{21},X_{22}]=X_{21}\,.$
Since (2.9) is in the Itô form, expression (2.5)) shows that it can be related to a Stratonovich operator of the form
$\mathfrak{S}(t,W_{1},x,y)=\left(\begin{pmatrix}0&1\\ -\omega_{f}^{2}&-k-\sigma^{2}/2\end{pmatrix}\begin{pmatrix}x\\ y\end{pmatrix},\begin{pmatrix}0&0\\ 0&-\sigma\end{pmatrix}\begin{pmatrix}x\\ y\end{pmatrix}\right).$
Hence, one can write
$\mathfrak{S}(t,W_{1},x,y)=\left(X_{12}-\omega_{f}^{2}X_{21}-(k+\sigma^{2}/2)X_{22},-\sigma X_{22}\right),$
which turns (2.9) into a stochastic Lie system and $V_{D}$ into an associated Vessiot–Guldberg Lie algebra. Note that more general models can be considered by assuming the coefficients $\omega_{f}$, $\sigma$, or $k$ depending on $t$ and/or $W_{1}$. One can also assume $\mathfrak{S}_{1},\mathfrak{S}_{2}$ to have as components more general linear combinations of the elements of $V_{D}$ with coefficients depending on $W_{1}$ and $t$. These models will still be stochastic Lie systems. $\triangle$
Example 2.4.
It is worth noting that an Itô stochastic differential equation may seem to have the form (2.8) without being a stochastic Lie system. This shows that the stochastic case does not matches exactly the form given in classical Lie systems in the Itô realm. Let us provide an example of this from the SIS model in Example 2.1. Let us study a stochastic differential equation that is related to the deterministic SIS model for particular values $N=100$, $\beta=1/2$, $\mu=\gamma=0$, which is indeed a deterministic approximation of it. Consider the Itô stochastic differential equation
$\delta I=I(50-I/2)\delta t+I\sigma(100-I)\delta\mathcal{B}\,,$ (2.11)
for a certain $t$-dependent parameter $\sigma=\sigma(t)$, which generalises the stochastic SIS system studied in [32, p. 880]. Consider the vector fields
$Y_{1}=\frac{\partial}{\partial I},\qquad Y_{2}=I\frac{\partial}{\partial I},\qquad Y_{3}=I^{2}\frac{\partial}{\partial I},$
which span a Lie algebra of vector fields with commutation constants
$[Y_{1},Y_{2}]=Y_{2},\qquad[Y_{1},Y_{3}]=2Y_{2},\qquad[Y_{2},Y_{3}]=Y_{3}.$
Moreover, the operator of the form
$\mathcal{I}(t,\mathcal{B},I)=(50Y_{2}-1/2Y_{3},100\sigma(t)Y_{2}-\sigma(t)Y_{3}),$
is related to the Vessiot–Guldberg Lie algebra $\langle Y_{1},Y_{2},Y_{3}\rangle$ and one would expected (2.11) to be a stochastic Lie system. Nevertheless, one has that (2.11) is related to the Stratonovich differential equation
$\delta I=\left(-\sigma(t)I^{3}+\left(-\frac{1}{2}+150\sigma^{2}(t)\right)I^{2}+I(50-5000\sigma^{2}(t))\right)\delta t+I\sigma(t)(100-I)\circ\delta\mathcal{B}\,,$ (2.12)
which cannot be described by the Vessiot–Guldberg Lie algebra $\langle Y_{1},Y_{2},Y_{3}\rangle$ and any other larger one (all of them are a generic point isomorphic on $\mathbb{R}$ to a Lie subalgebra of $\mathfrak{sl}_{2}$ (see [20, 30] for details). Indeed, one can prove, as shown hereafter in Section 3, that (2.12) is not a stochastic Lie system. $\triangle$
There are Itô stochastic differential equations whose coefficients match the form of the coefficients in (2.8) and they are still stochastic Lie systems. This is due to the fact that the transformation (2.5) maps the initial (2.2) into a new stochastic differential equation that retains again the condition (2.8). Notwithstanding, this is not the general case.
The following result will be of utility for applications.
Definition 2.5.
We call $t$-dependent stochastic Riccati differential equation the Stratonovich stochastic differential equation on $\mathbb{R}$ with stochastic variables given by the semi-martingales $\mathcal{B}^{2},\ldots,\mathcal{B}^{\ell}:\mathbb{R}_{+}\times\Omega\rightarrow\mathbb{R}$ of the form
$\delta\Gamma=\left(\sum_{\alpha=0}^{2}b_{\alpha}(t)\Gamma^{\alpha}\right)\delta t+\sum_{\beta=2}^{\ell}\sum_{\alpha=0}^{2}b_{\beta\alpha}(t)\Gamma^{\alpha}\circ\delta\mathcal{B}^{\beta},\qquad\forall\Gamma\in\mathbb{R},\qquad\forall t\in\mathbb{R}.$ (2.13)
for arbitrary $t$-dependent functions $b_{\alpha}(t),b_{\mu\alpha}(t)$ with $\alpha=0,1,2$ and $\mu=2,\ldots,\ell$. Observe that $\Gamma^{\alpha}$ here denotes the stochastic variable $\Gamma$ raised to the $\alpha$-th power.
$t$-Dependent stochastic Riccati differential equations are stochastic Lie systems. The following proposition is immediate.
Proposition 2.6.
An Itô differential equation of the form
$\delta\Gamma=\left(\sum_{\alpha=0}^{2}b_{\alpha}(t)\Gamma^{\alpha}\right)\delta t+\sum_{\beta=2}^{\ell}\sum_{\alpha=0}^{1}b_{\beta\alpha}(t)\Gamma^{\alpha}\delta\mathcal{B}^{\beta},\qquad\forall\Gamma\in\mathbb{R},\qquad\forall t\in\mathbb{R}$
for arbitrary $t$-dependent functions $b_{\alpha}(t),b_{\mu\alpha}(t)$ with $\alpha=0,1,2$ and $\mu=2,\ldots,\ell$ is also a noise-free stochastic Riccati differential equation.
It is worth noting that the theory of Lie systems can be generalised to many different realms [20, 23]. In particular, there is the theory of foliated Lie systems [17]. This suggests us the following generalisation.
Definition 2.7 (Foliated stochastic Lie system).
A stochastic foliated Lie system is a stochastic system of differential equations on $M$ of the form
$\delta\Gamma=\mathfrak{S}(B,\Gamma)\circ\delta B\,,$ (2.14)
where $B:\mathbb{R}_{+}\times\Omega\rightarrow\mathbb{R}^{\ell}$ is an $\mathbb{R}^{\ell}$-valued semi-martingale and $\mathfrak{S}(B,\Gamma):\mathrm{T}_{B}\mathbb{R}^{\ell}\to\mathrm{T}_{\Gamma}M$ is a Stratonovich operator, such that
$\mathfrak{S}_{j}(B,\Gamma)=\sum_{\alpha=1}^{r}b_{j}^{\alpha}(B,\Gamma)Y_{\alpha}(\Gamma)\,,\qquad j=1,\ldots,\ell\,,\qquad\forall B\in\mathbb{R}^{\ell},\qquad\forall\Gamma\in M,$ (2.15)
and the vector fields $\{Y_{1},\ldots,Y_{r}\}$ on $M$ span an $r$-dimensional Vessiot–Guldberg Lie algebra such that $b^{\alpha}_{j}(B,\Gamma)$, with $\alpha=1,\ldots,r$ and $j=1,\ldots,\ell$, are first integrals of the vector fields $Y_{1},\ldots,Y_{r}$ for every $B\in\mathbb{R}^{\ell}$.
It is worth noting that there are many potential applications concerning the generalisation to the stochastic realm of famous types of Lie systems such as matrix, projective Riccati equations, Bernouilli equations, and so on. As illustrated in this section, stochastic Lie systems have applications in predator-prey models, oscillator type models, et cetera [3, 35, 52]. Moreover, nonlinear stochastic differential equations are difficult to study. Notwithstanding, under certain conditions, its linearisation can describe its equilibrium properties [3]. Linear or even affine stochastic differential equations with stochastic variables related to semi-martingales are stochastic Lie systems (cf. [39]).
3 Superposition rules and stochastic Lie systems
Let us study the superposition rule notion for stochastic differential equations and the characterisation of stochastic differential equations admitting a superposition rule. This will lead to review and slightly correct some mistakes in the main theorem in [39].
Definition 3.1.
A superposition rule for a Stratonovich stochastic differential equation of the form (2.3) on a manifold $M$ is a function $\Phi:M^{m+1}\to M$ such that, for a generic set $\Gamma_{1},\ldots,\Gamma_{m}:\mathbb{R}_{+}\times\Omega\rightarrow M$, of particular solutions of (2.3), the general solution $\Gamma$ to (2.3) reads
$\Gamma=\Phi(z;\Gamma_{1},\ldots,\Gamma_{m}),$
where $z\in M$ is a point to be related to initial conditions.
It is remarkable that superposition rules for stochastic differential equations do not depend on $\mathbb{R}^{\ell}$. Let us introduce now several concepts that will be useful to describe and calculate superposition rules for stochastic Lie systems.
The diagonal prolongation to $M^{k}$ of a vector bundle $\tau:F\to M$ is the vector bundle $\tau^{[k]}:F^{k}=F\times\overset{(k)}{\dotsb}\times F\mapsto M^{k}=M\times\overset{(k)}{\dotsb}\times M$, of the form
$\tau^{[k]}(f_{(1)},\dotsc,f_{(k)})=(\tau(f_{(1)}),\dotsc,\tau(f_{(k)}))\,,\qquad\forall f_{(1)},\ldots,f_{(k)}\in F\,,$
with fibers of the form
$F^{k}_{(x_{(1)},\ldots,x_{(k)})}=F_{x_{(1)}}\oplus\cdots\oplus F_{x_{(k)}}\,,\qquad\forall(x_{(1)},\ldots,x_{(k)})\in M^{k}\,.$ (3.1)
Every section $e:M\to F$ of the vector bundle $\tau$ has a natural diagonal prolongation to a section $e^{[k]}$ of the vector bundle $\tau^{[k]}$ given by
$e^{[k]}(x_{(1)},\ldots,x_{(k)})=(e(x_{(1)}),\cdots,e(x_{(k)}))\,,\qquad\forall(x_{(1)},\ldots,x_{(k)})\in M^{k}.$
If we consider that every $e(x_{(a)})$ takes values in the $a$-th copy of $F$ within (3.1), one can write
$e^{[k]}(x_{(1)},\ldots,x_{(k)})=e(x_{(1)})+\cdots+e(x_{(k)})\,,\qquad\forall(x_{(1)},\ldots,x_{(k)})\in M^{k},$
which is a simple useful notation for applications. The diagonal prolongation of a function $f\in\mathscr{C}^{\infty}(M)$ to $M^{k}$ is the function on $M^{k}$ given by
$f^{[k]}(x_{(1)},\ldots,x_{(k)})=f(x_{(1)})+\ldots+f(x_{(k)})\,.$
Consider also the sections $e^{(j)}$ of $\tau^{[k]}$, where $j\in\{1,\ldots,k\}$ and $e$ is a section of $\tau$, given by
$e^{(j)}(x_{(1)},\dots,x_{(k)})=0+\cdots+e(x_{(j)})+\cdots+0\,,\qquad\forall(x_{(1)},\ldots,x_{(k)})\in M^{k}\,.$ (3.2)
If $\{e_{1},\ldots,e_{r}\}$ is a basis of local sections of the vector bundle $\tau$, then $e_{i}^{(j)}$, with $j=1,\dotsc,k$ and $i=1,\dotsc,r$, is a basis of local sections of $\tau^{[k]}$. For simplicity, we will frequently write $e(x_{(j)})$ instead of $e^{(j)}$ if it is clear what we mean.
Due to the obvious canonical isomorphisms
$(\mathrm{T}M)^{[k]}\simeq\mathrm{T}M^{k}\quad\text{and}\quad(\mathrm{T}^{\ast}M)^{[k]}\simeq\mathrm{T}^{\ast}M^{k}\,,$
the diagonal prolongation $X^{[k]}$ of a vector field $X\in\mathfrak{X}(M)$ can be understood as a vector field ${X}^{[k]}$ on $M^{k}$, and the diagonal prolongation, $\alpha^{[k]}$, of a one-form $\alpha$ on $M$ can be understood as a one-form ${\alpha}^{[k]}$ on $M^{k}$.
More explicitly, if $Y$ be a vector field on a manifold $M$. The diagonal prolongation of $Y$ to $M^{k}$ is the vector field
${Y}^{[k]}(x_{(1)},\ldots,x_{(k)})=\sum_{a=1}^{k}Y(x_{(a)})$
on $M^{k}$ obtained by considering $\mathrm{T}M^{k}\simeq\mathrm{T}M\times\dotsb\times\mathrm{T}M$ ($k$ times). Even more particularly, if $Y=x\frac{\partial}{\partial x}$ is a vector field on $\mathbb{R}$, then $Y^{[k]}=\sum_{a=1}^{k}x_{(a)}\frac{\partial}{\partial x_{(a)}}$.
The diagonal prolongation of a vector field can be extended to $t$-dependent vector fields on $M$, namely mappings $X\colon\mathbb{R}\times M\rightarrow\mathrm{T}M$ such that $X(t,\cdot)$ is a standard vector field $M$, by assuming that the diagonal extension to $M^{k}$ of the $t$-dependent vector field $X$ on $M$ is the $t$-dependent vector field $\widetilde{X}^{[k]}$ on $M^{k}$ whose value for every fixed valued of $t$, let us say $\widetilde{X}^{[k]}_{t}$, is the diagonal prolongation to $M^{k}$ of the vector field $X_{t}$. The space of diagonal prolongations of vector fields in $\mathfrak{X}(M)$ to $M^{k}$ form a Lie subalgebra of $\mathfrak{X}(M^{k})$. In fact, the mapping $X\in\mathfrak{X}(M)\mapsto{X}^{[k]}\in\mathfrak{X}(M^{k})$ is a Lie algebra morphism, i.e. it is a linear mapping such that
$[Y_{1},Y_{2}]^{[k]}=\left[{Y}^{[k]}_{1},{Y}^{[k]}_{2}\right],\qquad\forall Y_{1},Y_{2}\in\mathfrak{X}(M)\,.$ (3.3)
Let us solve a mistake in the proof of the stochastic Lie theorem in [39]. The issue appears in the direct part of the statement [39, p. 215, Theorem 3.1]. In particular, this makes [39, Remark 3.3.(i)] incorrect. For the sake of completeness, we will state that part of the work adopting our notation and writing in full the contents referenced by the labels used in [39]. We hereafter refer to
“Moreover, the fact
$\left[{Z}^{[k+1]}_{1},(Z_{2}+\lambda Z_{3})^{[k+1]}\right]=([Z_{1},Z_{2}+\lambda Z_{3}])^{[k+1]},\qquad\forall Z_{1},Z_{2},Z_{3}\in\mathfrak{X}(M),\qquad\lambda\in\mathbb{R},$
and the hypothesis on $\{Y_{1},\dotsc,Y_{\kappa}\}$ being in involution [^2] It is really meant that the distribution spanned by $Y_{1},\ldots,Y_{\kappa}$ is involutive and $Y_{1},\ldots,Y_{\kappa}$ are linearly independent over $\mathbb{R}$. imply by the classical Frobenius Theorem that
$D=\left\langle{Y}^{[k+1]}_{1},\ldots,{Y}^{[k+1]}_{\kappa}\right\rangle$
is integrable.”
which is false. Let us explain the mistake and build a counterexample. Our counterexample will be relevant as it will show that certain stochastic differential equations in the Itô approach are not stochastic Lie systems and do not accept a superposition rule.
The problem relies on the fact that if $Y_{1},\dotsc,Y_{\kappa}$ span an involutive distribution on $M$, then the diagonal prolongations $\widetilde{Y}^{[k+1]}_{1},\ldots,\widetilde{Y}^{[k+1]}_{\kappa}$ do not span the distribution given by the diagonal prolongations of the distribution spanned by $Y_{1},\ldots,Y_{\kappa}$.
For instance, consider the two vector fields on =$\mathbb{R}_{\circ}=\mathbb{R}\setminus\{0\}$ given by
$Y_{1}=x^{2}\frac{\partial}{\partial x}\,,\qquad Y_{2}=x^{3}\frac{\partial}{\partial x}$
that span an involutive distribution on =$\mathbb{R}_{\circ}=\{x\in\mathbb{R}:x\neq 0\}$. Notwithstanding, their diagonal prolongations, ${Y}^{[s]}_{1},{Y}^{[s]}_{2}$, to $\mathbb{R}_{\circ}^{s}$ , with $s>2$, do not need to span an involutive distribution. We hereafter consider $s>2$. Indeed, one has the diagonal prolongations
${Y}_{1}^{[s]}=\sum_{a=1}^{s}x_{(a)}^{2}\frac{\partial}{\partial x_{(a)}},\qquad{Y}_{2}^{[s]}=\sum_{a=1}^{s}x_{(a)}^{3}\frac{\partial}{\partial x_{(a)}}$
on $\mathbb{R}^{s}_{\circ}$. Their successive commutators become, as stated in [39], diagonal prolongations of an element of the involutive distribution spanned by $Y_{1},Y_{2}$, namely $\mathrm{T}\mathbb{R}_{\circ}$. In particular,
$\operatorname{ad}_{{Y}^{[s]}_{1}}^{k}{Y}^{[s]}_{2}=\stackrel{{\scriptstyle k-{\rm times}}}{{\overbrace{\left[{Y}_{1}^{[s]},[\ldots,[\ldots,[\ldots,[{Y}_{1}^{[s]}\right.}}}\left.,{Y}_{2}^{[s]}]\ldots]\ldots]\ldots]\right]=\sum_{a=1}^{s}k!x_{(a)}^{3+k}\frac{\partial}{\partial x_{(a)}}=\left(k!x^{k+3}\frac{\partial}{\partial x}\right)^{[s]},$
for $k=1,2,\ldots$ Notwithstanding, any $\operatorname{ad}_{{Y}^{[s]}_{1}}^{k}{Y}^{[s]}_{2}$, with $k\in\{2,3,4,\ldots\}$, does not take values in the distribution $\mathcal{D}=\left\langle{Y}^{[s]}_{1},{Y}^{[s]}_{2}\right\rangle$ at a generic point of $\mathbb{R}_{\circ}^{s}$. Even worse, the smallest (in the sense of inclusion) involutive generalised distribution containing ${Y}^{[s]}_{1},{Y}^{[s]}_{2}$ spans the whole tangent space to $\mathbb{R}_{\circ}^{s}$ at almost every point and a superposition rule for a system described by a generic combination of $Y_{1},Y_{2}$ does not exist as it must be constructed from the non-constant first integrals of the vector fields taking values in an involutive distribution containing ${Y}^{[s]}_{1},{Y}^{[s]}_{2}$ for some $s>2$ (see the proof for the Lie theorem in [18, 20, 23, 39]). In fact, the vector fields ${Y}^{[s]}_{\mu}=\sum_{a=1}^{s}x_{(a)}^{\mu+1}\frac{\partial}{\partial x_{(a)}}$, with $\mu=1,\ldots,s$, on $\mathbb{R}_{\circ}^{s}$ are linearly independent almost everywhere. To verify this fact, it is enough to see that the determinant of their coefficients in the basis $\frac{\partial}{\partial x_{(a)}}$, with $a=1,\ldots,s$, read
$\begin{vmatrix}x_{(1)}^{2}&x_{(2)}^{2}&\cdots&x_{(s)}^{2}\\ x_{(1)}^{3}&x_{(2)}^{3}&\cdots&x_{(s)}^{3}\\ \vdots&\vdots&&\vdots\\ x_{(1)}^{s+1}&x_{(2)}^{s+1}&\cdots&x_{(s)}^{s+1}\end{vmatrix}=x_{(1)}^{2}\dotsb x_{(s)}^{2}\begin{vmatrix}1&1&\cdots&1\\ x_{(1)}&x_{(2)}&\cdots&x_{(s)}\\ \vdots&\vdots&&\vdots\\ x_{(1)}^{s-1}&x_{(2)}^{s-1}&\cdots&x_{(s)}^{s-1}\end{vmatrix}=\prod_{a=1}^{s}x_{(a)}^{2}\prod_{1\leq i<j\leq s}(x_{(j)}-x_{(i)})\,.$
which makes that the smallest involutive generalised distribution containing ${Y}^{[s]}_{1},{Y}^{[s]}_{2}$ on $\mathbb{R}_{\circ}^{s}$ to be equal to $\mathrm{T}\mathbb{R}_{\circ}^{s}$ almost everywhere. This makes the existence of common non-constant first integrals for ${Y}^{[s]}_{1},{Y}^{[s]}_{2}$, which will be common non-contact first integrals for all $Y^{[s]}_{1},Y^{[s]}_{2},\ldots$ impossible.
The previous counterexample is very important as it concerns the stochastic generalisations of the so-called Abel equations (see [15] and references therein). Hence, the mistake in [39] has potential practical consequences. In particular, note that the SIS model in the Itô form (2.11), for a non-constant function $\sigma(t)$, takes the Stratonovich form (2.12). If one tries to write the $t$-dependent coefficient for $\delta t$ as a linear combination with $t$-dependent functions of a family of vector fields on $\mathbb{R}$ spanning a finite-dimensional Lie algebra, one finds that, for a generic $t$-dependent function $\sigma(t)$, one has to obtain a finite-dimensional Lie algebra of vector fields on $\mathbb{R}$ containing $x^{2}\frac{\partial}{\partial x},x^{3}\frac{\partial}{\partial x}$, which is impossible as shown above.
Despite the above mistake, the stochastic Lie–Scheffers theorem in [39] can be corrected by assuming that the initial family of vector field $Y_{1},\ldots,Y_{r}$ close an $r$-dimensional Lie algebra. Note that the assumption that $Y_{1},\ldots,Y_{r}$ are linearly independent over $\mathbb{R}$ is necessary. Otherwise, e.g. if $Y_{r-1}=Y_{r}$, one obtains that the vector fields $Y_{1}^{[m]},\ldots,Y_{r}^{[m]}$ are always linearly dependent and Lemma 3.2 of [39] cannot be applied in [39, p. 918]. On the other hand, if $Y_{1},\ldots,Y_{r}$ are linearly independent, their diagonal prolongations become linearly independent at a generic point exactly when $m$ is such that they span a distribution of rank $r$ at a generic point (cf. [20]). Then, Lemma 3.2 can be applied normally as accomplished in [39, p. 918].
Theorem 3.2 (Stochastic Lie theorem).
Let
$\delta\Gamma=\mathfrak{S}(B,\Gamma)\circ\delta B$ (3.4)
be a stochastic differential equation on $M$, where $B:\mathbb{R}_{+}\times\Omega\rightarrow\mathbb{R}^{\ell}$ is a given $\mathbb{R}^{\ell}$-valued semi-martingale and $\mathfrak{S}(B,\Gamma):\mathrm{T}_{B}\mathbb{R}^{\ell}\to\mathrm{T}_{\Gamma}M$, for every $B\in\mathbb{R}^{\ell}$ and $\Gamma\in M$, is a Stratonovich operator from $\mathbb{R}^{\ell}$ to $M$. Then, (3.4) admits a superposition rule if and only if
$\mathfrak{S}_{j}(B,\Gamma)=\sum_{\alpha=1}^{r}b_{j}^{\alpha}(B)Y_{\alpha}(\Gamma)\,,\qquad j=1,\ldots,\ell\,,$ (3.5)
for every $B\in\mathbb{R}^{\ell}$ and $\Gamma\in M$, where the vector fields $\{Y_{1},\ldots,Y_{r}\}$ span an $r$-dimensional Vessiot–Guldberg Lie algebra on $M$.
Proof.
Let us prove the part of the direct implication of the theorem in [39, Theorem 3.1] that requires some comments due to our correction. One has that there always exists a number $m$ such that the diagonal prolongation of a basis $Y_{1},\ldots,Y_{r}$ of the Vessiot–Guldberg Lie algebra of the Lie system reaches rank $r$ at a generic point. It was proved in [20] that this happens exactly when the diagonal prolongations span a distribution of rank $r$ almost everywhere. Let be $\mathcal{D}_{0}$ the distribution spanned by vector fields $Y^{[m+1]}_{1},\ldots,Y^{[m+1]}_{r}$. Then $\mathcal{D}_{0}$ has constant rank on an open subset of $M^{m+1}$. Since the Lie brackets $[Y_{i},Y_{j}]$ are linear combinations of the vector fields of $V$, with constant coefficients, then their diagonal prolongations to $M^{m+1}$ are linear combinations, with constant coefficients, of the $Y^{[m+1]}_{1},\ldots,Y^{[m+1]}_{r}$ and $\mathcal{D}_{0}$ is integrable. Then, one can extend $\mathcal{D}_{0}$ to an integrable distribution $\mathcal{D}$ of corank $n$ in $M^{m+1}$, i.e. $\mathcal{D}_{0}\subset\mathcal{D}$ on every point of an open neighbourhood of a point in $M^{m+1}$. This can always be done in such a manner that the leaves of $\mathcal{D}$ project diffeomorphically onto an open subset of $M^{m}$ via the canonical projection $\pi:(\Gamma_{(1)},\ldots,\Gamma_{(m+1)})\in M^{m+1}\mapsto(\Gamma_{(1)},\ldots,\Gamma_{(m)})\in M^{m}$. The obtained expression is a local superposition rule for (3.4) as shown in [39, Theorem 3.1] in basis of [39, Proposition 2.4]. ∎
4 Hamiltonian stochastic Lie systems
This section studies stochastic Lie systems admitting a Vessiot–Guldberg Lie algebra of Hamiltonian vector fields relative to a geometric structure. In particular, we mainly focus on Hamiltonian vector fields relative to a symplectic form, but our analysis can be immediately extended to more general Hamiltonian systems, e.g. related to Poisson manifolds. To motivate our approach, let us consider a stochastic differential equation on $\mathbb{R}^{2}$ of the form
$\begin{cases}\dfrac{dx}{dt}=x({a}_{1}(t,\mathcal{B})-b_{1}y),\\ \dfrac{dy}{dt}=y({a}_{1}(t,\mathcal{B})+b_{2}x),\end{cases}$ (4.1)
for a function $a_{1}\in\mathscr{C}^{\infty}(\mathbb{R}^{2})$ and constants $b_{1},b_{2}\in\mathbb{R}$. This system is a stochastic generalisation of a type of Palomba–Goodwin model, which in turn is a particular case of Lotka–Volterra system [31, 49]. Our stochastic generalisation is obtained by assuming that the standard $t$-dependent coefficients depend, additionally, on a semi-martingale $\mathcal{B}$. In this case, the associated Stratonovich operator takes the form
$\mathfrak{S}(t,\mathcal{B},x,y)=\begin{pmatrix}x({a}_{1}(t,\mathcal{B})-b_{1}y)\\ y({a}_{1}(t,\mathcal{B})+b_{2}x)\end{pmatrix}\,,$
which allows us to write
$\delta\Gamma=\begin{pmatrix}\delta x\\ \delta y\end{pmatrix}=\mathfrak{S}(t,\mathcal{B},x,y)\delta t\,.$
Moreover, $\mathfrak{S}(t,\mathcal{B},x,y)$ can be written as
$\mathfrak{S}(t,\mathcal{B},x,y)=a_{1}(t,\mathcal{B})X_{1}+X_{2}\,,$
where the vector fields $X_{1}$ and $X_{2}$ are given by
$X_{1}=x\frac{\partial}{\partial x}+y\frac{\partial}{\partial y}\,,\qquad X_{2}=xy\left(-b_{1}\frac{\partial}{\partial x}+b_{2}\frac{\partial}{\partial y}\right).$
Since $[X_{1},X_{2}]=-X_{2},$ the vector fields $X_{1},X_{2}$ span a non-Abelian two-dimensional Lie algebra $\mathfrak{h}_{2}$. Hence, system (4.1) is a stochastic Lie system.
The vector fields $X_{1},X_{2}$ are Hamiltonian relative to the symplectic structure $\omega=\frac{1}{xy}\mathrm{d}x\wedge\mathrm{d}y$ on $\mathcal{O}=\{(x,y)\in\mathbb{R}^{2}:xy\neq 0\}$. In fact,
$\iota_{X_{1}}\omega=\mathrm{d}\ln|y/x|\,,\qquad\iota_{X_{2}}\omega=\mathrm{d}\left(-b_{1}y-b_{2}x\right).$
Hence, we call (4.1), a Hamiltonian stochastic Lie system on $\mathcal{O}$. Moreover, (4.1) could be further generalised, e.g. by considering
$\delta\Gamma=\mathfrak{S}(t,\mathcal{B},x,y)\delta t+\mathfrak{S}^{\prime}(t,\mathcal{B},x,y)\circ\delta\mathcal{B},$
where $\mathfrak{S}^{\prime}(t,\mathcal{B},x,y)=f_{1}(t,\mathcal{B})X_{1}+f_{2}(t,\mathcal{B})X_{2}$ for arbitrary functions $f_{1},f_{2}\in\mathscr{C}^{\infty}(\mathbb{R}^{2})$.
Let us analyse the stochastic SIS model in the Stratonovich approach given by [22]
$\delta S$ $=\left[\Lambda-\mu_{1}S-\frac{\beta SI}{\kappa+S}+\gamma I\right]\delta t-\sigma S\circ\delta\mathcal{B}\,,$
$\delta I$ $=\left[\frac{\beta SI}{\kappa+S}-(\mu_{1}+\gamma)I\right]\delta t-\sigma I\circ\delta\mathcal{B}\,,$
where $\kappa$ is a rate of disease-related death, $\Lambda$ is an input of new members, $\mu$ is a natural mortality, $\beta$ is a transmission rate, $\gamma$ is a rate of recovery, $\sigma^{2}$ is the intensity of white noise, $\mu_{1}=\mu+\sigma^{2}/2$, and $\mathcal{B}$ is a Brownian motion. Let us consider the case $\kappa=0$, namely
$\delta S$ $=\left[\Lambda-\mu_{1}S+(\gamma-\beta)I\right]\delta t-\sigma S\circ\delta\mathcal{B}$
$\delta I$ $=\left[-\mu_{1}I-(\gamma-\beta)I\right]\delta t-\sigma I\circ\delta\mathcal{B}$ (4.2)
one can consider the vector fields on $\mathbb{R}^{2}$ given by
$\bar{X}_{1}=S\frac{\partial}{\partial S}+I\frac{\partial}{\partial I}\,,\quad\bar{X}_{2}=I\frac{\partial}{\partial S}-I\frac{\partial}{\partial I}\,,\quad\bar{X}_{3}=\frac{\partial}{\partial S}.$
Then,
$[\bar{X}_{1},\bar{X}_{3}]=-\bar{X}_{1}\,,\qquad[\bar{X}_{2},\bar{X}_{3}]=0,\qquad[\bar{X}_{1},\bar{X}_{2}]=0.$
Hence, the system becomes a stochastic Lie system. In fact, the Stratonovich operator reads
$\mathfrak{S}(B,\Gamma)=(\Lambda\bar{X}_{3}-\mu_{1}\bar{X}_{1}+(\gamma-\beta)\bar{X}_{2},-\sigma\bar{X}_{1})\,,$ (4.3)
where we consider solutions $\Gamma:(t,\omega_{p})\in\mathbb{R}_{+}\times\Omega\mapsto(S,I)\in\mathcal{U}=\{(S,I)\in\mathbb{R}^{2}:SI\neq 0\}$. In the particular case of $\Lambda=0$, one can consider the locally defined symplectic form, away from $(0,0)$, of the form
$\omega=\mathrm{d}\xi_{1}\wedge\mathrm{d}\xi_{2}$
where $\xi_{1},\xi_{2}$ are local coordinates such that $\bar{X}_{1}=\frac{\partial}{\partial\xi_{1}}$ and $\bar{X}_{2}=\frac{\partial}{\partial\xi_{2}}$. The existence of $\xi_{1},\xi_{2}$ is due to the fact that $\bar{X}_{1}\wedge\bar{X}_{2}$ does not vanish and $[\bar{X}_{1},\bar{X}_{2}]=0$ away from $(0,0)$. Note that system (4.2) can be generalised to consider $t$-dependent coefficients. the stochastic differential equation on $\mathbb{R}^{n}$ with a Wiener process $W_{1}$ (see [35, 36, 57] for analogues on $\mathbb{R}$) in Stratonovich form given by
$\delta\begin{pmatrix}x_{i}\\ y_{i}\end{pmatrix}=\begin{pmatrix}\lambda&a\\ b&-\lambda\end{pmatrix}\begin{pmatrix}x_{i}\\ y_{i}\end{pmatrix}\delta t+\begin{pmatrix}\lambda^{\prime}&a^{\prime}\\ b^{\prime}&-\lambda^{\prime}\end{pmatrix}\begin{pmatrix}x_{i}\\ y_{i}\end{pmatrix}\circ\delta W_{1}\,,\qquad i=1,\ldots,n\,,$ (4.4)
which is defined on $\mathrm{T}^{\ast}\mathbb{R}^{n}$. This recovers Example 2.3, namely Eq. (2.9), by taking $k=0$ and $\sigma=0$ as a special case. It can be proved that (2.9) is not Hamiltonian relative to any symplectic form on $\mathbb{R}^{2}$ for arbitrary $t$-dependent coefficients as it leads to a Vessiot–Guldberg Lie algebra on the plane related to $\mathfrak{gl}_{2}$ [23]. But (4.4) can now be related to a Vessiot–Guldberg Lie algebra $V$ on $\mathrm{T}^{*}\mathbb{R}^{n}$ spanned by
$X_{1}=\sum_{i=1}^{n}y_{i}\frac{\partial}{\partial x_{i}}\,,\qquad X_{2}=\frac{1}{2}\sum_{i=1}^{n}\left(x_{i}\frac{\partial}{\partial x_{i}}-y_{i}\frac{\partial}{\partial y_{i}}\right)\,,\qquad X_{3}=-\sum_{i=1}^{n}x_{i}\frac{\partial}{\partial y_{i}}\,.$
Since,
$[X_{1},X_{2}]=X_{1}\,,\qquad[X_{1},X_{3}]=X_{2}\,,\qquad[X_{2},X_{3}]=X_{3}\,,$
this Lie algebra is isomorphic to $\mathfrak{sl}_{2}$ and it is the diagonal prolongation to $(\mathrm{T}^{*}\mathbb{R})^{n}$ of a Vessiot–Guldberg Lie algebra isomorphic to $\mathfrak{sl}_{2}$ on $\mathrm{T}^{*}\mathbb{R}$ (see [6, 23]). Moreover, it is known to be a Lie algebra of Hamiltonian vector fields relative to the symplectic form
$\omega=\sum_{i=1}^{n}\mathrm{d}x_{i}\wedge\mathrm{d}y_{i}\,.$
Hence, $X_{1},X_{2},X_{3}$ have Hamiltonian functions given by
$h_{1}=\sum_{i=1}^{n}\frac{y_{i}^{2}}{2}\,,\qquad h_{2}=\frac{1}{2}\sum_{i=1}^{n}x_{i}y_{i}\,,\qquad h_{3}=\sum_{i=1}^{n}\frac{x_{i}^{2}}{2}\,.$
relative to $\omega$ that span a Lie algebra isomorphic to $\mathfrak{sl}_{2}$.
Now, (4.4) can be related to Stratonovich operator of the form
$\mathfrak{S}(x,y,t,W)=\left(\lambda X_{2}+aX_{1}-bX_{3},\lambda^{\prime}X_{2}+a^{\prime}X_{1}-b^{\prime}X_{3}\right),$
which turns (4.4) into a Hamiltonian stochastic Lie system and $V$ into an associated Vessiot–Guldberg Lie algebra isomorphic to $\mathfrak{sl}_{2}$.
Definition 4.1.
A Hamiltonian stochastic Lie system on a manifold $M$ relative to a probability space $\Omega$ and a semi-martingale $B\colon\mathbb{R}_{+}\times\Omega\rightarrow\mathbb{R}^{\ell}$ is a stochastic Lie system related to a Stratonovich operator $\mathfrak{H}$ admitting an associated Vessiot–Guldberg Lie algebra on $M$ consisting of Hamiltonian vector fields relative to some geometric structure on $M$.
Of course, the above definition means that there are Hamiltonian stochastic Lie systems related to symplectic, Poisson, Jacobi, contact or Dirac geometries, among others. It is worth noting that we are mainly interested in Hamiltonian stochastic Lie systems given by
$\delta\Gamma=\sum_{\alpha=1}^{r}b_{1}^{\alpha}(t)X_{\alpha}\delta t+\sum_{i=2}^{\ell}\sum_{\alpha=1}^{r}b_{i}^{\alpha}(t)X_{\alpha}\circ\delta\mathcal{B}^{i}$ (4.5)
such that the vector fields $X_{1},\ldots,X_{r}$ span an $r$-dimensional Vessiot–Guldberg Lie algebra of Hamiltonian vector fields on $M$ relative to a symplectic form. Note that the coefficients in (4.5) depend only on time, although dependence on $\mathcal{B}^{2},\dotsc,\mathcal{B}^{\ell}$ will be also analysed. In our main case of study, (4.5) is then related to an $\ell$-family of $t$-dependent Hamiltonian functions being each of them a linear combination with coefficients depending on $t$ of certain Hamiltonian functions $h_{1},\dotsc,h_{r}\in\mathscr{C}^{\infty}(M)$ contained in a finite-dimensional Lie algebra of Hamiltonian functions relative to the Poisson bracket of the symplectic manifold. It may happen that the Hamiltonian functions $h_{1},\ldots,h_{r}$ need to be enlarged with an additional constant function to close on a Lie algebra, but this option is hereafter skipped for simplicity. In particular, the $\ell$-family of $t$-dependent vector fields in (4.5) have $t$-dependent Hamiltonian functions
$\widetilde{h}_{1}=\sum_{\alpha=1}^{r}b^{\alpha}_{1}(t)h_{\alpha}\,,\qquad\widetilde{h}_{j}=\sum_{\alpha=1}^{r}b^{\alpha}_{j}(t)h_{\alpha}\,,\qquad j=2,\ldots,\ell\,,$
respectively. It is standard to call $\widetilde{h}_{1}$ the Hamiltonian of the stochastic differential equation.
Figure 1: These are two representations of the evolution in terms of the time of a particular solution to the Hamiltonian Lie system $\delta y=-x\delta t-Ax\delta W,\delta x=y\delta t+Ay\delta W$, with initial condition (.$(.2,0)$ a semi-martingale $W$, and a parameter $A$. The symplectic structure is $\omega=\mathrm{d}x\wedge\mathrm{d}y$. The system has a strong constant of motion $x^{2}+y^{2}$. The constant of the motion is always conserved, but solutions jump back and forward relative to the deterministic solution.
5 Stability for stochastic Hamiltonian systems
Let us recall the stability theory for stochastic differential equations and apply it to stochastic Hamiltonian systems. In particular, we will be specially interested in Hamiltonian stochastic Lie systems. For a survey on the theory for general stability of stochastic differential equations and other related results see [4, 40, 52].
Consider a stochastic differential equation on $M$ of the particular form
$\delta\Gamma=\mathfrak{S}_{t}(t,\Gamma)\delta t+\mathfrak{S}_{\mathcal{B}}(t,\Gamma)\circ\delta\mathcal{B}\,,\qquad\forall\Gamma\in M,\qquad\forall t\in\mathbb{R},$ (5.1)
where $B=(t,\mathcal{B})\colon\mathbb{R}_{+}\times\Omega\rightarrow\mathbb{R}^{\ell}$ is an $\mathbb{R}^{\ell}$-valued semi-martingale and
$\mathfrak{S}(t,\Gamma)=(\mathfrak{S}_{t}(t,\Gamma),\mathfrak{S}_{\mathcal{B}}(t,\Gamma))\colon\mathrm{T}_{B}\mathbb{R}^{\ell}\to\mathrm{T}_{\Gamma}M$
is the associated Stratonovich operator. As in previous sections, it is assumed that initial conditions are values of $M$ which are taken with probability equal to one. Note that coefficients of the Stratonovich operator are to depend only on $t$ and $\Gamma$. Stochastic differential equations of this type are common in the literature [11].
An equilibrium point for the stochastic differential equation (5.1) is a point $\Gamma_{e}\in M$ such that
$\mathfrak{S}(t,\Gamma_{e})=0\,,\qquad\forall t\in\mathbb{R}\,.$
In this case, the stochastic disturbance, which is described by $\mathfrak{S}_{\mathcal{B}}$, does not act at the equilibrium point $\Gamma_{e}$. Note that if $\mathfrak{S}_{\mathcal{B}}$ is not assumed to be zero at the equilibrium point for every $t\in\mathbb{R}$, solutions may move away from the point depending on the values of the stochastic variable (see Figure 2).
Figure 2: Representation of a solution to $\delta x=y\delta t-0.008\delta W,\delta y=-x\delta t-0.008\delta W$ on $\mathbb{R}^{2}$ with initial condition $(0,0)$ and stochastic variable given by a semi-martingale $W$, showing that a non-vanishing stochastic part of the stochastic differential equation may move solutions away from a deterministic equilibrium point.
Definition 5.1.
Consider the Stratonovich differential equation
$\delta\Gamma=\mathfrak{S}(t,\Gamma)\circ\delta B$ (5.2)
with solutions $\Gamma\colon\mathbb{R}_{+}\times\Omega\rightarrow M$. Given $\Gamma_{0}\in M$ and $s\in\mathbb{R}$, denote by $\Gamma^{s,\Gamma_{0}}$ the particular solution of (5.2) such that $\Gamma^{s,\Gamma_{0}}(\omega_{0})=\Gamma_{0}$ for all $\omega_{0}\in\Omega$. Let $\Gamma_{e}\in M$ be an equilibrium point of (5.2). The equilibrium point $\Gamma_{e}$ is almost surely (Lyapunov) stable if for any open neighbourhood $U$ of $\Gamma_{e}$ there exists a neighbourhood $\mathcal{U}$ of $\Gamma_{e}$ such that for any $\Gamma^{\prime}\in\mathcal{U}$, one has that $\Gamma^{s,\Gamma^{\prime}}\subset U$ almost surely (a.s.).
Let us provide a relevant analogue of the method used in Hamiltonian symplectic systems to verify stability [1]. But before that, let us provide a definition for strongly conserved quantities for Hamiltonian stochastic differential equations (see [40, Definition 2.2]).
Definition 5.2.
A function $f\in\mathscr{C}^{\infty}(M)$ is said to be strongly conserved of a stochastic Hamiltonian (5.2) if, for any particular solution $\Gamma$ with initial condition $\Gamma_{0}$, we have that $f(\Gamma)=f(\Gamma_{0})$.
Note that strongly conserved quantities can also be defined for systems (5.2) whose Stratonovich operator depend also on $B$. The most relevant result for our purposes in the study of stability of Hamiltonian stochastic Lie systems is the following proposition (see [40, Theorem 2.15] and references therein).
Proposition 5.3.
(Stochastic Dirichlet’s Criterion) Assume that there exists a function $f\in\mathscr{C}^{\infty}(M)$ so that $\mathrm{d}f_{\Gamma_{e}}=0$ and that the quadratic form $\mathrm{d}^{2}f_{\Gamma_{e}}$ is positive or negative definite. If $f$ is a strongly conserved quantity on the solutions of (5.2), then the equilibrium point $\Gamma_{e}$ is almost surely stable.
Let us turn now to Hamiltonian stochastic differential equations on symplectic manifolds. In this case, we focus on stochastic differential equations of the type
$\delta\Gamma=\mathfrak{H}(t,\Gamma)\circ\delta B\,,$ (5.3)
where $B\colon\mathbb{R}_{+}\times\Omega\rightarrow\mathbb{R}^{\ell}$ is a semi-martingale and $\mathfrak{H}:\mathrm{T}\mathbb{R}^{\ell}\times M\rightarrow\mathrm{T}M$ is a Stratonovich operator such that
$\mathfrak{H}(t,\Gamma)=(Y_{1}(t,\Gamma),\ldots,Y_{\ell}(t,\Gamma))\,,$
where $Y_{1},\ldots,Y_{\ell}$ stand for $t$-dependent Hamiltonian vector fields relative to a symplectic form $\omega$ on $M$ with $t$-dependent Hamiltonian functions $\widetilde{h}_{1},\ldots,\widetilde{h}_{\ell}\in\mathscr{C}^{\infty}(\mathbb{R}\times M)$, respectively. The symplectic structure in this case allows for the study of the system via the Hamiltonian functions of the system, which gives powerful methods to study their properties. Note that (5.3) has indeed an associated function $\widetilde{h}\colon\mathbb{R}\times M\rightarrow\mathbb{R}^{\ell}$ of the form
$\widetilde{h}=(\widetilde{h}_{1},\ldots,\widetilde{h}_{\ell}).$
Recall that $\widetilde{h}_{1}$ is normally called the Hamiltonian of the system, but Hamiltonian stochastic differential equations have several associated Hamiltonians. In fact, note that $\widetilde{h}_{1}$ need not be conserved even when it is $t$-independent. It is immediate that it is not a constant of the motion in general as $\{\widetilde{h}_{1},\widetilde{h}_{i}\}$ not need to be zero in general. Notwithstanding, it is remarkable that a $f\in\mathscr{C}^{\infty}(M)$ is a strong constant of motion for (4.5) if
$\{(\widetilde{h}_{\alpha})_{t},f\}=0\,,\qquad\alpha=1,\ldots,\ell\,,\qquad\forall t\in\mathbb{R}.$
Let us assume that there exists a function $f$ that is a constant of motion of $\delta\Gamma$ and it also satisfies that it has a minimum at an equilibrium point. Then, it is immediate that $\delta\Gamma$ has a stable equilibrium point.
It is simple to apply the above results to linear or affine stochastic differential equations depending on a semi-martingale $B$, which are stochastic Lie systems, and to choose cases with a Vessiot–Guldberg Lie algebra of Hamiltonian vector fields relative to a symplectic structure.
6 Relative equilibrium points and stochastic Hamiltonian Lie systems
In short, relative equilibrium points for a system of differential equations on a manifold $M$, in general, and for a system of stochastic differential ones, in particular, are points where the dynamics is generated by symmetries of the differential equations. In this section, we define relative equilibrium points for Hamiltonian stochastic differential equations in a Stratonovich form of the particular type (5.3) using the theory in [38, 40].
Let us start by reviewing the notion of symmetries for stochastic differential equations and symplectic reduction for Hamiltonian stochastic differential equations.
Definition 6.1.
Let $B:\mathbb{R}_{+}\times\Omega\rightarrow\mathbb{R}^{\ell}$ be a semi-martingale and let $\mathfrak{S}\colon\mathrm{T}\mathbb{R}^{\ell}\times M\rightarrow\mathrm{T}M$ be a Stratonovich operator. A diffeomorphism $\phi\colon M\rightarrow M$ is a symmetry of the stochastic differential equation associated with $\mathfrak{S}$ if
$\mathfrak{S}(\delta B,\phi(\Gamma))=\mathrm{T}\phi[\mathfrak{S}(\delta B,\Gamma)]\,,\qquad\forall(\delta B,\Gamma)\in\mathrm{T}\mathbb{R}^{\ell}\times M\,.$
More generally, a Lie group action $\Phi:G\times M\rightarrow M$ is a symmetry of the stochastic differential equation induced by $\mathfrak{S}$ if $\Phi_{g}:\Gamma\in M\mapsto\Phi(g,\Gamma)\in M$ is a symmetry of $\mathfrak{S}$ for every $g\in G$. Similarly, we say that $\phi$ and $\Phi$ are symmetries of the Stratonovich operator $\mathfrak{S}$.
Remark 6.2.
Note the slight change of notation for the Stratonovich operator in Definition 6.1, where the first entry is an element of $\mathrm{T}\mathbb{R}^{\ell}$ instead of an element of $\mathbb{R}^{\ell}$ as in previous sections.
The relevance of symmetries of Stratonovich operators is due to the fact that they transform a particular solution of the stochastic differential equations related to it to another particular solution of the same stochastic differential equation [38, Theorem 2.2].
Let us give a $t$-dependent generalisation of the stochastic Marsden–Meyer–Weinstein reduction for Hamiltonian stochastic systems as introduced in [41, Theorem 3.1] for the case of Hamiltonian stochastic systems (5.3), which essentially amounts to applying a standard reduction for each particular $t\in\mathbb{R}$.
Let $(M,\omega)$ be a symplectic manifold and let $\Phi\colon G\times M\rightarrow M$ be a Lie group action admitting a coadjoint equivariant momentum map $J\colon M\rightarrow\mathfrak{g}^{*}$ being also a Lie group symmetry of the Hamiltonian functions $\widetilde{h}\colon\mathbb{R}\times M\rightarrow\mathbb{R}^{\ell}$ of a Stratonovich operator $\mathfrak{H}$. Every regular $\mu\in\mathfrak{g}^{*}$ of $J$ gives rise to a function $h_{\mu}\colon\mathbb{R}\times J^{-1}(\mu)/G_{\mu}\rightarrow\mathbb{R}^{\ell}$ on the manifold $M_{\mu}=J^{-1}(\mu)/G_{\mu}$ determined by the equality $h_{\mu}(t,\pi_{\mu}(\cdot))=h(t,\iota_{\mu}(\cdot))$, where $\iota_{\mu}\colon J^{-1}(\mu)\hookrightarrow M$ is the natural immersion and we assume $G_{\mu}$ to act freely and properly on $J^{-1}(\mu)$. In turn, this induces a stochastic Hamiltonian system on the symplectic reduced space $(M_{\mu}:=J^{-1}(\mu)/G_{\mu},\omega_{\mu})$ whose Stratonovich operator $\mathfrak{H}_{\mu}\colon\mathrm{T}\mathbb{R}^{\ell}\times M_{\mu}\to\mathrm{T}M_{\mu}$ is given by (c.f. [38, Theorem 3.1])
$\mathfrak{H}_{\mu}(\delta B,[\Gamma]))=\mathrm{T}_{\Gamma}\pi_{\mu}(\mathfrak{H}(\delta B,\Gamma))\,,\qquad\forall\delta B\in\mathrm{T}\mathbb{R}^{\ell}\,,\qquad\forall\Gamma\in M\,.$
Note that $\mathrm{T}_{\Gamma}\pi_{\mu}$ is assumed to act on each component of $\mathfrak{H}(t,\Gamma)\colon\mathrm{T}_{t}\mathbb{R}^{\ell}\rightarrow\mathrm{T}_{\Gamma}M$ for each $t\in\mathbb{R}$ in the natural manner. Moreover, if $\Gamma$ is a solution of the stochastic Hamiltonian system associated with $\mathfrak{H}$ with initial condition $\Gamma_{0}\subset J^{-1}(\mu)$ at $t=0$, then so is $\Gamma_{\mu}:=\pi_{\mu}(\Gamma)$ with respect to $\mathfrak{H}_{\mu}$, with initial condition $\pi_{\mu}(\Gamma_{0})$ at $t=0$.
In the particular case of stochastic Hamiltonian Lie systems, the reduced stochastic equation is a stochastic Hamiltonian Lie system related to a Stratonovich operator of the form
$\mathfrak{H}_{\mu}(t,[\Gamma]))=\left(\sum_{\alpha=1}^{r}b^{1}_{\alpha}(t)X_{h^{\mu}_{\alpha}}([\Gamma]),\ldots,\sum_{\alpha=1}^{r}b^{\ell}_{\alpha}(t)X_{h^{\mu}_{\alpha}}([\Gamma])\right),$
where the $h^{\mu}_{\alpha}\in\mathscr{C}^{\infty}(M_{\mu})$ are determined by the conditions $h^{\mu}_{\alpha}\circ\pi_{\mu}=h_{\alpha}\circ\iota_{\mu}$ from the generators of the Lie algebra of Hamiltonian functions $h_{1},\ldots,h_{r}$ of the initial Hamiltonian stochastic Lie system. Note that in the case of Hamiltonian stochastic Lie systems with a Vessiot–Guldberg Lie algebra $V$, we require all related Hamiltonian functions, $h_{1},\ldots,h_{r}$ to be invariant relative to the action of the Lie group $\Phi\colon G\times M\rightarrow M$.
The above discussion suggests us the following definition.
Definition 6.3.
Given a Hamiltonian stochastic differential equation (5.3) on $M$, a relative equilibrium point $\Gamma_{\rm rel}\in M$ of (5.3) is a point such that
$\mathfrak{H}_{i}(t,\Gamma_{\rm rel})\in D_{\Gamma_{\rm rel}},\qquad i=1,\ldots,\ell\,,\qquad\forall t\in\mathbb{R}\,,$ (6.1)
where $D$ is the distribution generated by the fundamental vector fields of a Lie group action of symmetries $\Phi\colon G\times M\rightarrow M$ with a momentum map $J\colon M\rightarrow\mathfrak{g}^{*}$.
To understand the above condition in terms of the better known characterisation for standard Hamiltonian systems of differential equations [46], consider a basis $\{\xi_{M}^{1},\ldots,\xi^{r}_{M}\}$ of fundamental vector fields of the Lie group action $\Phi$ corresponding to a basis $\{\xi_{1},\ldots,\xi_{r}$} of $\mathfrak{g}$. Then, the relation (6.1) amounts to saying that
$\mathfrak{H}(t,\Gamma_{\rm rel})=\left(\sum_{\alpha=1}^{r}f_{1}^{\alpha}(t)\xi^{\alpha}_{M}(\Gamma_{\rm rel}),\ldots,\sum_{\alpha=1}^{r}f_{\ell}^{\alpha}(t)\xi^{\alpha}_{M}(\Gamma_{\rm rel})\right)\,,\qquad\forall t\in\mathbb{R},$
for certain $t$-dependent functions $f_{i}^{\alpha}\in\mathscr{C}^{\infty}(\mathbb{R})$ for $i=1,\ldots,\ell$ and $\alpha=1,\ldots,r$.
Let $\mu=J(\Gamma_{\rm rel})$ in the above definition. If the Hamiltonian stochastic differential equation (5.3) can be reduced, then $\Gamma_{\rm rel}\in M$ is contained in $J^{-1}(\mu)$. If it is assumed to be a submanifold of $M$, the evolution of $\mathfrak{H}$ is contained within $J^{-1}(\mu)$. This implies that the right-hand side of (6.1) belongs indeed to $D_{\Gamma_{\rm rel}}\cap\mathrm{T}_{\Gamma_{\rm rel}}J^{-1}(\mu)$ Moreover, the restriction of $\mathfrak{H}$ to it can be reduced to $J^{-1}(\mu)/G_{\mu}$. Moreover, the $\mathfrak{H}_{\mu}(t,\pi_{\mu}(J^{-1}(\mu)))=0$. In other words, one has the following proposition, whose inverse is analogue to its deterministic case.
Proposition 6.4.
Let $\Gamma_{\rm rel}\in M$ be a relative equilibrium point for a stochastic Hamiltonian system (5.3) and let $J(\Gamma_{\rm rel})=\mu$ be a regular point of $J$, let $G_{\mu}$ act on $J^{-1}(\mu)$ freely and properly so that $J^{-1}(\mu)/G_{\mu}$ becomes a manifold. Then, the reduction of (5.3) to $M_{\mu}=J^{-1}(\mu)/G_{\mu}$ is such that $\pi_{\mu}(\Gamma_{\rm rel})$ is an equilibrium point of the reduced system and vice versa.
Finally, let us characterise relative equilibrium points for stochastic Hamiltonian systems.
Theorem 6.5.
(Stochastic Relative Equilibrium Theorem) A point $\Gamma_{\rm rel}\in M$ is a relative equilibrium point for a Hamiltonian stochastic differential equation (4.5) related to $\widetilde{h}:\mathbb{R}\times M\rightarrow\mathbb{R}^{\ell}$ if and only if there exist elements $\xi_{t}=(\xi^{1}_{t},\ldots,\xi_{t}^{\ell})\in\mathfrak{g}^{\ell}$ for $t\in\mathbb{R}$ such that $(\xi_{t},\Gamma_{\rm rel})$ are critical points of the coordinates of the functions $\widehat{h}_{t}:\mathfrak{g}^{\ell}\times M\rightarrow\mathbb{R}^{\ell}$, given by
$\widehat{h}_{t}^{\alpha}(\xi_{t},\Gamma)=\widetilde{h}^{\alpha}(t,\Gamma)-\left\langle{J}(\Gamma)-\mu_{e},\ \xi_{t}^{\alpha}\right\rangle,\qquad\alpha=1,\ldots,\ell,$
for $\mu_{e}:=J(\Gamma_{e})$.
Proof.
Let us prove the direct part. At the critical points $(\xi_{t},\Gamma_{\rm rel})$, the functions $\widehat{h}_{t}(\xi_{t},\cdot):M\rightarrow\mathbb{R}^{\ell}$ have a critical point $\Gamma_{\rm rel}\in M$. Hence, each coordinate of $\widehat{h}_{t}(\xi_{t},\cdot)$ satisfies that $\mathrm{d}(\widetilde{h}_{t}^{\alpha}-\langle J(\cdot),\xi^{\alpha}_{t}\rangle)|_{\Gamma_{\rm rel}}=0$. This implies that $(X_{\widetilde{h}_{t}^{\alpha}})|_{\Gamma_{\rm rel}}=X_{J_{\xi^{\alpha}_{t}}}|_{\Gamma_{\rm rel}}$ and the point $\Gamma_{\rm rel}$ is a relative equilibrium point of $\mathfrak{H}$. Conversely, if $\Gamma_{\rm rel}$ is a relative equilibrium point to $\mathfrak{H}$, at each coordinate of $\mathfrak{H}$, one has that $(X_{\widetilde{h}^{\alpha}})_{t}(\Gamma_{\rm rel})=(\xi_{t})_{M}^{\alpha}(\Gamma_{\rm rel})$ for certain $\xi_{t}^{1},\ldots,\xi_{t}^{\ell}\in\mathfrak{g}$ with $t\in\mathbb{R}$. Consequently, one has that $X_{\widetilde{h}_{t}^{\alpha}-\langle J,\xi^{\alpha}_{t}\rangle}=0$. This implies that each one of these vector fields has a Hamiltonian function given by $\widetilde{h}_{t}^{\alpha}-\langle J-\mu_{e},\xi_{t}^{\alpha}\rangle$ which has a critical point at the given point $\Gamma_{\rm rel}$. ∎
7 Stochastic Poisson coalgebra method
Let us review the Poisson coalgebra method for Hamiltonian systems relative to different geometric structures. In general, our approach slightly matches what can be found in previous works, but it also introduces certain simplifications and assumptions that allow one to apply it to many topics. In particular, the method can still be applied to Hamiltonian stochastic Lie systems by considering that they are determined by a certain family of Hamiltonian functions. Indeed, our present procedure is a new modification of the coalgebra method for deriving superposition rules for $k$-symplectic Lie systems devised in [24].
It is convenient to stress that the proof of the stochastic Lie theorem shows that a superposition rule for stochastic Lie systems, in Stratonovich form, can be obtained in a similar manner to the case of deterministic Lie systems, namely (see [18, 20, 23] for details and examples):
  1. 1.
    Consider a Vessiot–Guldberg Lie algebra spanned by a basis of vector fields $X_{1},\ldots,X_{r}$ on the manifold $M$ where the stochastic Lie system is defined on.
  2. 2.
    Find the smallest natural number $m\in\mathbb{N}$, so that $X^{[m]}_{1},\ldots,X_{r}^{[m]}$ are linearly independent at a generic point.
  3. 3.
    Use local coordinates $x^{1},\ldots,x^{n}$ on $M$ and consider this coordinate system to be defined on each copy of $M$ within $M^{m+1}$ to get a coordinate system $\{x^{i}_{(a)}\mid i=1,\ldots,n,\,\,a=0,\ldots,m\}$ on $M^{m+1}$. Obtain $n$ functionally independent first-integrals $F_{1},\ldots,F_{n}$ common to all diagonal prolongations $X^{[m+1]}_{1},\ldots,X^{[m+1]}_{r}$ such that
    $\frac{\partial(F_{1},\ldots,F_{n})}{\partial(x_{(0)}^{1},\ldots,x_{(0)}^{n})}\neq 0.$ (7.1)
  4. 4.
    The condition (7.1) allows us to ensure that the equations $F_{i}=k_{i}$, for $i=1,\ldots,n$, enable us to write the expressions of the variables $x_{(0)}^{1}\ldots,x_{(0)}^{n}$ in terms of $x^{1}_{(a)},\ldots,x_{(a)}^{n}$, with $a=1,\ldots,m,$ and $k_{1},\ldots,k_{n}$.
  5. 5.
    The obtained expressions lead to a superposition rule depending on a generic family of $m$ particular solutions and the constants $k_{1},\ldots,k_{n}$. This holds true even in the Stratonovich stochastic realm.
Note that every Hamiltonian stochastic Lie system is related to a Stratonovich operator $\mathfrak{H}$ given by $\ell$ components and each one can be understood as a $t$-dependent vector field (that can be also $\mathbb{R}^{\ell}$-dependent if we allow the dependence on all stochastic variables). Hence, $\mathfrak{H}$ can be understood as an $t$-dependent $\ell$-vector field, i.e. for every $t\in\mathbb{R}$, it is a section of the $\ell$-tangent bundle of velocities $\mathrm{T}^{\ell}M=\mathrm{T}M\oplus\dotsb\oplus\mathrm{T}M\rightarrow M$, where $\mathrm{T}M\oplus\dotsb\oplus\mathrm{T}M$ is to be considered as the Whitney sum of the tangent bundle to $M$ with itself. In this manner, one can prolong it to a section $\mathfrak{H}$ of ×$\mathbb{R}\times(\mathrm{T}^{k}M)^{[m]}\rightarrow\mathbb{R}\times M^{m}$. Moreover, $\mathfrak{H}$ is related to a certain $t$-dependent Hamiltonian function $h\colon\mathbb{R}\times M\rightarrow\mathbb{R}^{\ell}$. Then, $h$ is a section of ×$\mathbb{R}\times M\times\mathbb{R}^{\ell}\rightarrow\mathbb{R}\times M$ and it can be extended to a section of ×$\mathbb{R}\times M^{m}\times\mathbb{R}^{\ell}\rightarrow\mathbb{R}\times M^{m}$. Moreover, the prolongation $h^{[m]}$ is the $t$-dependent Hamiltonian function related to $\mathfrak{H}^{[m]}$. Then, the following result is immediate.
Proposition 7.1.
If $\mathfrak{H}$ is a Hamiltonian stochastic Lie system admitting a $t$-dependent Hamiltonian $h\colon\mathbb{R}\times M\rightarrow\mathbb{R}^{\ell}$ relative to a symplectic form $\omega$, then $\mathfrak{H}^{[m]}$ is a Hamiltonian stochastic Lie system relative to the symplectic form $\omega^{[m]}$ admitting a Hamiltonian $h^{[m]}\colon\mathbb{R}\times M^{m}\rightarrow\mathbb{R}^{\ell}$. In particular, if $h_{1},\ldots,h_{r}$ is a basis of the Lie algebra of Hamilton functions for $\mathfrak{H}$, then $h_{1}^{[m]},\ldots,h_{r}^{[m]}$ is a basis of the Lie algebra of Hamiltonian functions for $\mathfrak{H}^{[m]}$.
One has the following immediate result.
Proposition 7.2.
The space of $\ell$-Hamiltonian functions on a symplectic manifold $(M,\omega)$, with Poisson bracket {,}$\{\cdot,\cdot\}_{\omega}$ induced by $\omega$, is a Poisson algebra relative to the bracket
$\{h,h^{\prime}\}_{\ell}=(\{h_{1},h^{\prime}_{1}\}_{\omega},\ldots,\{h_{\ell},h^{\prime}_{\ell}\}_{\omega})\,,\qquad\forall h=(h_{1},\ldots,h_{\ell}),h^{\prime}=(h^{\prime}_{1},\ldots,h^{\prime}_{\ell})\in\mathscr{C}^{\infty}(M,\mathbb{R}^{\ell})\,,$
and the multiplication
$h\cdot h^{\prime}=(h_{1}h_{1}^{\prime},\ldots,h_{\ell}h^{\prime}_{\ell})\,,\qquad\forall h,h^{\prime}\in\mathscr{C}^{\infty}(M,\mathbb{R}^{\ell})\,.$
It is worth recalling that the space of $\ell$-Hamiltonian functions relative to an $\ell$-symplectic structure was not a Poisson algebra due to the fact that the multiplication of $\ell$-Hamiltonian functions could not be ensured to be an $\ell$-Hamiltonian function as each coordinate is assumed to be a Hamiltonian function of the same vector field and this cannot be ensured for their bracket as each coordinate of the bracket is associated with different presymplectic forms and may be related to different vector fields [24]. Here, this problem does not appear, since each $h_{i}$, with $i=1,\ldots,\ell$, may be the Hamiltonian function of a different vector field.
For the sake of completeness, let us prove the following result.
Proposition 7.3.
Let $\mathfrak{H}$ be a Hamiltonian stochastic Lie system with respect to a symplectic form $\omega$ and possessing a smallest (in the sense of inclusion) Lie–Hamiltonian Vessiot–Guldberg Lie algebra $(\mathfrak{W},\{h_{1},\ldots,h_{r}\})$ relative to the Poisson bracket related to $\omega$. A function $f\in\mathscr{C}^{\infty}(M)$ is a strong constant of the motion for $\mathfrak{H}$ if commutes with all the elements of ${\rm Lie}(\{h_{t}\}_{t\in\mathbb{R}},\{\cdot,\cdot\})$. In particular, $f$ is a strong contact of the motion if it commutes with the elements of $\mathfrak{M}$.
Proof.
The function $f$ is a strong constant of the motion for $\mathfrak{H}$ if
$0=X^{i}_{t}f\,,\qquad\forall t\in\mathbb{R}\,,\qquad i=1,\ldots,\ell,$ (7.2)
where the $X_{t}^{i}$ are the components of the $t$-dependent $\ell$-vector field related to $\mathfrak{H}$. Since $X_{t}^{i}=\{h^{i}_{t},f\}$ for $i=1,\ldots,\ell$, the result follows. Note also that
$\{f,\{h_{t},h_{t^{\prime}}\}_{\omega}\}_{\omega}=\{\{f,h_{t}\},h_{t^{\prime}}\}_{\omega}\}_{\omega}+\{h_{t},\{f,h_{t^{\prime}}\}_{\omega}\}_{\omega}\,,\qquad\forall t,t^{\prime}\in\mathbb{R}\,.$
Inductively, $f$ is shown to commute with all the elements of ${\rm Lie}(\{h_{t}\}_{t\in\mathbb{R}})$. In particular, ${\rm Lie}(\{h_{t}\}_{t\in\mathbb{R}})\subset\mathfrak{W}$ ensures that $f$ is a strong constant of motion.
In Lie–Hamilton systems, a $t$-dependent Hamiltonian vector field admitted for every $t\in\mathbb{R}$ a Hamiltonian function belonging to a finite-dimensional Lie algebra of Hamiltonian functions relative to a Poisson bracket.
Proposition 7.4.
Let $\mathfrak{H}$ be a Hamiltonian stochastic Lie system with an associated Lie–Hamilton Vessiot–Guldberg Lie algebra $(\mathfrak{W}=\langle h_{1},\ldots,h_{r}\rangle,\{\cdot,\cdot\}_{\omega})$ relative to the symplectic form $\omega$. Let $\{v_{1},\ldots,v_{r}\}$ be a basis of linear coordinates on $\mathfrak{g}^{*}\simeq\mathfrak{W}$. Given the momentum map $J:M\to\mathfrak{g}^{\ast}$, the pull-back $J^{\ast}C$ of any Casimir function $C$ on $\mathfrak{g}^{\ast}$ is a constant of the motion for $\mathfrak{H}$. Moreover, if $h_{i}=J^{*}v^{i}$ for $i=1,\ldots,r$, and $C=C(v_{1},\ldots,v_{r})$, then
$C\left(\sum_{a=1}^{m}h_{1}(x_{(a)}),\ldots,\sum_{a=1}^{m}h_{r}(x_{(a)})\right)\,,$ (7.3)
is a constant of the motion of $\mathfrak{H}^{[m]}$.
The stochastic Poisson coalgebra method takes its name from the fact that it is applied to Hamiltonian stochastic Lie systems and it analyses the use of Poisson coalgebras and a so-called coproduct to obtain superposition rules. In fact, the coproduct is responsible for the form of (7.3) (see [7]).
8 Conclusions and outlook
This paper has reviewed in a simple manner the theory of stochastic Lie systems solving a mistake in the literature. Some relations between the Stratonovich and the Itô approaches to stochastic Lie systems have been reviewed, which has shown that the stochastic Lie theorem in the Itô approach has a different non-expected form. Types of Hamiltonian stochastic Lie systems have been introduced as well as some possible generalisations. We have also introduced the theory of Hamilton stochastic Lie systems. The theory of stability and relative equilibrium points for Hamiltonian stochastic differential equations has been developed and, in particular, we have focused in the study of Hamiltonian stochastic Lie systems relative to symplectic forms. Some results concerning the energy-momentum method for stochastic Lie systems have been obtained, generalizing [5]. Many examples concerning physical/epidemiological applications have been developed. We have also remarked that the theory of Poisson coalgebras can be developed for our Hamiltonian stochastic Lie systems.
Concerning applications, stochastic SIS models have been analysed. Such models use to be studied using deterministic methods in the literature. Instead, we here approach them in a pure stochastic approach. Many other physical systems have also been analysed.
In the future, we aim to further develop the energy-momentum method for stochastic Lie systems. This entails the search for criteria ensuring our of a Hamiltonian Stratonovich operator the stability of equilibrium points obtained after reduction. Moreover, we expect, as in the deterministic case, to obtain certain degeneracies that will have to be analysed to solve the problem. Note that a further study of the advantages of epidemiological models from a pure stochastic point of view will have to be addressed. Moreover, we also plan to study a stochastic Hamiltonian system appearing in celestial mechanics concerning the stochastic variation of the inertia tensor [52]. Finally, it would be interesting to study other types of superposition rules appearing in stochastic differential equations [37] and applying the Poisson coalgebra method for Hamiltonian stochastic Lie systems .
Acknowledgements
Fruitful discussions with J.P. Ortega about setting a correction for his stochastic Lie theorem are acknowledged. X. Rivas acknowledges partial financial support from the Spanish Ministry of Science and Innovation grants PID2021-125515NB-C21 and RED2022-134301-T of AEI, and the Ministry of Research and Universities of the Catalan Government project 2021 SGR 00603. J. de Lucas acknowledges a MINIATURA-5 grant by the Polish Narodowe Centrum Nauki under project Nr 2021/05/X/ST1/01797. The research of Marcin Zając was funded by the Polish National Science Centre under the contract number UMO-2021/41/N/ST1/02908. For the purpose of Open Access, the author has applied a CC-BY public copyright licence to any Author Accepted Manuscript (AAM) version arising from this submission. The authors also acknowledge the contribution of RED2022-134301-T funded by MCIN/AEI/10.13039/501100011033 (Spain).